Find f ◦ g and g ◦ f , where f (x) = 3x - 2 and g(x) = 2x + 2, are functions from R to R

1
Expert's answer
2022-04-21T13:32:34-0400

"(f\circ g)(x)=f(g(x))=\\
3g(x)-2=3(2x+2)-2=6x+4"


"(g\circ f)(x)=g(f(x))=\\
=2f(x)+2=2(3x-2)+2=6x-2"